OpenStudy (anonymous):

A long distance phone company has a monthly fee of $7.95 and charges a rate of $0.05 per minute. Another long distance company has a monthly fee of $9.95 and charges a rate of $0.03 per minute. At how many minutes would the two companies have equal charges?

OpenStudy (anonymous):

10 minutes 50 minutes 32 minutes 100 minutes

OpenStudy (anonymous):

100 minutes

OpenStudy (anonymous):

100 Minutes

OpenStudy (anonymous):

Thanks for the help :D

OpenStudy (anonymous):

x*0,05+7,95=x*0,03+9,95 x=100